# Entropy and multiplicity of hadrons in the high energy limit within dipole cascade models

### v1

@@ I Introduction
If the hadron–parton duality [33] is true then the final state multiplicities should also be maximal entropy distributions. Recent work showed that the observed distributions in proton-proton collisions are compatible with the idea of the application of the maximal entropy principle [34, 35] in the final state [36]. It was also observed that the shape of the final state multiplicity distributions is different if they are measured in disjoint, equally wide pseudorapidity intervals (e.g. LHCb [37] and H1 [18]) or in symmetric windows, centered to mid-rapidity with increasing width (e.g. ALICE [38, 39], ATLAS [40, 41, 42], CMS [43], UA5 [44]). However, while the distributions can be derived from the principle of maximum entropy, the comparison with data measured using different rapidity-range definitions requires extra attention.
@@ II.2 Generalized Mueller dipole model
which is a generalization of Eq. (2) with an extra parameter $h$. The equation has been obtained in the context of study of Krylov complexity [52, 53, 54] for the coherent states of the SL(2,R) group and mapped to QCD in Ref. [47]. Here $h$, is a conformal weight of SL(2,R) and for BFKL case $h=1/2$.
In our setup, the parameter $h$ measures the deviation from the 1D Mueller dipole model. While the Eq. (9) can be written in such form that after setting $h=1/2$ it reduces to Eq. (2) in our studies, we use a version of this model first introduced in Ref. [47], since in this formulation one can also account for $p_{0}$, which effectively corresponds to the vacuum contribution. Such contributions were accounted for in recent studies [19, 51] and allowed for successful description of hadronic entropy data.
@@ III Data and fit description
The analyzed data sets are described in Tab. 1. However, the psuedorapidity variable is not defined uniquely, by introducing the variable $\ln\langle n\rangle$, we are able to compare the obtained entropy among experiments. The details of the effect of the rapidity definition are discussed in Ref. [36] and in Appendix A. A comprehensive overview of multiplicity distributions and their shape can be found in Ref. [59].
@@ Acknowledgments
S.L. is grateful for the support of the Miniatura 7 grant of the National Science Centre of Poland. S.L. thanks the hospitality and insightful discussions to M. Płoskoń in Berkeley and G. Contreras in Prague. The authors would like to thank the fruitful discussions to P. Caputa, M. Hentschinski, H. Jung, M. Praszałowicz, M. Šumbera, Z. Tu, D. Kharzeev and J. Otwinowski.
@@
Appendix A The maximal entropy principle
As it was shown in Refs. [71, 36] that the maximum entropy principle is a suitable framework to derive the functional form of the measured charged particle multiplicity distributions. We generalize the results that has been shown in Ref. [36] to describe data that follows negative binomial distribution with non-integer $k$ parameter.
The main assumption is that the particles are emitted in a random, memoryless process that can be modeled as a Poisson process. The mean of the Poisson process is not known but can be drawn from a continuous, maximal entropy distribution with fixed $\langle n\rangle$ and $\langle\ln n\rangle$: the Gamma distribution. Consequently, to determine the charged particle multiplicity distribution, the following integral should be calculated
$P(n)=$ $\int_{0}^{\infty}f(n|\mu)g(\mu)d\mu=\int_{0}^{\infty}\frac{\mu^{n}e^{-\mu}}{n!}\frac{\mu^{k-1}e^{\mu/\vartheta}}{\Gamma(k)\vartheta^{k}}d\mu=$ (13)
$=$ $\frac{\Gamma(k+n)}{\Gamma(k)\Gamma(n+1)}\left(\frac{1}{1+\vartheta}\right)^{k}\left(\frac{\vartheta}{1+\vartheta}\right)^{n}.$
This structure is usually referred to as the Poisson–Gamma mixture. The result of the integral is the negative binomial distribution [71, 36] that restores the geometric distribution as the $k=1$ special case.

### v2

@@ I Introduction
If the hadron–parton duality [33] is true then the final state multiplicities should also be maximal entropy distributions. Recent work showed that the observed distributions in proton-proton collisions are compatible with the idea of the application of the maximal entropy principle [34, 35, 36] in the final state [37]. It was also observed that the shape of the final state multiplicity distributions is different if they are measured in disjoint, equally wide pseudorapidity intervals (e.g. LHCb [38] and H1 [18]) or in symmetric windows, centered to mid-rapidity with increasing width (e.g. ALICE [39, 40], ATLAS [41, 42, 43], CMS [44], UA5 [45]). However, while the distributions can be derived from the principle of maximum entropy, the comparison with data measured using different rapidity-range definitions requires extra attention.
@@ II.2 Generalized Mueller dipole model
which is a generalization of Eq. (2) with an extra parameter $h$. The equation has been obtained in the context of study of Krylov complexity [53, 54, 55] for the coherent states of the conformal SL(2,R) group and mapped to QCD in Ref. [48]. Here $h$, is a conformal weight of this group which tels how the operators of the theory scale. This mapping is possible since the BFKL has been shown to obey conformal symmetry and in particular the conformal symmetry has been used to provide its solution in momentum [50] and impact parameter space [56]. In fact the leading eigenfunction of the BFKL kernel has conformal weight $h=1/2$. In our setup, the parameter $h$ measures the deviation from the 1D Mueller dipole model. While the Eq. (9) can be rewritten in such form that after setting $h=1/2$ and rescaling (see appendix A) it reduces to Eq. (2) in our studies, we use a version of this model first introduced in Ref. [48], since in this formulation one can also account for $p_{0}$, which effectively corresponds to the vacuum contribution. Such contributions were accounted for in recent studies [19, 52] and allowed for successful description of hadronic entropy data.
@@ III Data and fit description
The analyzed data sets are described in Tab. 1. However, the psuedorapidity variable is not defined uniquely, by introducing the variable $\ln\langle n\rangle$, we are able to compare the obtained entropy among experiments. The details of the effect of the rapidity definition are discussed in Ref. [37]. A comprehensive overview of multiplicity distributions and their shape can be found in Ref. [61].
@@ Acknowledgments
S.L. is grateful for the support of the Miniatura 7 grant of the National Science Centre of Poland and thanks the hospitality and insightful discussions to M. Płoskoń in Berkeley and G. Contreras in Prague. The authors would like to thank the fruitful discussions to P. Caputa, M. Hentschinski, H. Jung, M. Praszałowicz, M. Šumbera, Z. Tu, D. Kharzeev and J. Otwinowski.
@@
Appendix A Rescaling the general solution
In this appendix we show how the solution of the generalized model in Eq. (9) reduces to the Mueller model.
Let’s set $h=1/2$ and $C=1$ in the solution Eq. (10), that is then written as
$P_{n}(y)=e^{-\alpha y}(1-e^{-\alpha y})^{n}.$ (13)
Let’s multiply by $(1-e^{-\alpha y})$, so we obtain
$P_{n+1}(y)=e^{-\alpha y}(1-e^{-\alpha y})^{n+1}.$ (14)
Now, let’s multiply Eq. (9) by $(1-e^{-\alpha y})^{n+1}$, so we obtain
$\partial_{y}P_{n+1}(y)=\alpha nP_{n}(y)-\alpha(n+1)P_{n+1}(y)$ (15)
Substituting $n+1=k$, we obtain the Mueller model
$\partial_{y}P_{k}(y)=\alpha(k-1)P_{k-1}(y)-\alpha kP_{k}(y).$ (16)